OpenStudy (sk8teroy):

Find the real root of the equation 3z^3 - 10z^2 + 7z + 10 = 0 given that one root is 2 - i.

sam (.sam.):

\[3z^3 - 10z^2 + 7z + 10 = 0 \] \[(z+2-i) (3 z+2) (z+2+i) = 0\] \[3z+2=0\] \[z=-\frac{2}{3}\]

OpenStudy (sk8teroy):

How did you find the factor (3z+2) ?

OpenStudy (anonymous):

Multiply z+2-i with z+2+i and divide the product by the given equation to obtain 3z+2.

sam (.sam.):

@Aron West , The thing is he wants to know the roots of complex numbers first, Roots of complex numbers always conjugate to each other, so you'll get (z+2−i)=0 (z+2+i)=0 From 2 - i

OpenStudy (sk8teroy):

Thank you.

OpenStudy (anonymous):

Welcome!